### v1

# Many Worlds in Theory Space: A Quantum Origin for the Constants of Nature

Abstract
Modern physics relies on a specific set of dimensionless parameters—interaction strengths, mass ratios, and vacuum energy—whose observed values lie in extraordinarily narrow ranges required for complexity and life. Standard explanations often invoke external multiverses or ad hoc ensembles. This work proposes a purely quantum-mechanical origin for these constants by enlarging the configuration space of the Wheeler–DeWitt equation to include the theory-defining parameters themselves, and extends the Everett Many Worlds interpretation to include many worlds of different physical laws.
We treat fundamental constants not as fixed inputs but as dynamical quantum variables stabilized at the Grand Unified Theory (GUT) or Planck scale. The universal wavefunction thus carries amplitude over a “Theory Space” of possible physical laws. We demonstrate that early-universe symmetry breaking and decoherence naturally fragment this wavefunction into distinct, non-interacting “U-sectors,” each effectively a classical universe with fixed constants. In this extended Many-Worlds framework, fine-tuning is reinterpreted through quantum selection: while the vast majority of the cosmic wavefunction may describe sterile vacua, observers necessarily inhabit rare habitable branches. Finally, we suggest that the total integrated amplitude over theory space serves as a rigorous Bayesian evidence metric, enabling the comparative evaluation of candidate microscopic theories of quantum gravity.
1 Introduction
The “fine-tuning” of the universe refers to the empirical fact that the constants of nature, such as the fine-structure constant $\alpha$ or the cosmological constant $\Lambda$, occupy a tiny island of parameter space compatible with stable matter and long-lived stars [CarrRees1979, BarrowTipler1986].
The prevailing explanation for this puzzle relies on a synthesis of two conceptually distinct frameworks: the String Landscape and Eternal Inflation. It is useful to distinguish their roles:
  • The String Landscape serves as the “map” or the menu of possibilities. It is the static space of metastable vacua ($\sim 10^{500}$) derived from string theory, where each vacuum corresponds to a different compactification geometry and thus different effective laws of physics.
  • Eternal Inflation serves as the “engine.” It provides the dynamical mechanism to populate the landscape. Through continuous exponential expansion and bubble nucleation, it physically realizes the possibilities, creating vast, causally disconnected regions with different parameters.
Together, these form the Landscape Multiverse. Considerable theoretical effort has been directed toward understanding how these parameters could emerge from quantized degrees of freedom such as discrete fluxes and moduli fields. However, the Landscape Multiverse relies heavily on semiclassical reasoning; treating universes as bubbles nucleating in a background spacetime–and assumes the landscape structure a priori.
In contrast, the Many-Worlds Interpretation (MWI) of quantum mechanics [Everett1957] describes a multiverse generated internally by the unitary evolution of a single wavefunction. Standard MWI is limited, however: it generates branching histories (e.g., measurement outcomes) but not branching laws. All worlds in the standard MWI share the exact same electron mass and interaction strengths.
This paper unifies these perspectives by extending the canonical quantization of gravity. We propose that the “constants” of nature are quantum degrees of freedom that acquire fixed values through decoherence transitions in the very early universe. By enlarging the configuration space of the Wheeler–DeWitt equation [DeWitt1967], we construct a meta-wavefunction that superposes not just geometries, but entire theories.
In this framework, theory space ($\mathcal{T}$) acts as the landscape, but the engine that populates it is not inflationary nucleation, but quantum decoherence. The result is a quantum multiverse where the laws of physics themselves are subject to branching. The fine-tuning of our universe is thus reinterpreted as a quantum weighting problem: we inhabit a rare but habitable branch of the cosmic wavefunction.
The remainder of the paper proceeds as follows. Section 2 details the high-energy origin of the constants. Section 3 constructs the enlarged configuration space. Section 4 describes the quantum structure of constants. Section 5 presents the meta-Wheeler–DeWitt equation.
2 The Quantum Origin of Physical Constants
A foundational principle of quantum mechanics is that all physical observables must correspond to linear operators acting on a Hilbert space. Yet, in the standard formulation of quantum field theory and cosmology, the physical constants of the Standard Model–gauge couplings, masses, and mixing angles–are not represented as operators but appear merely as fixed c-numbers in the Lagrangian. This indicates a conceptual incompleteness: the theory specifies a quantum dynamics for geometry and matter fields while treating the theory-defining structure itself as rigid classical background data.
To resolve this, we elevate these parameters to dynamical quantum variables. In this framework, the “constants” we observe are not fundamental inputs but derivative quantities determined by the vacuum state of the high-energy theory. While phase transitions such as QCD confinement or Electroweak symmetry breaking occur at relatively low energies (post-inflation), the parameters governing these transitions (coupling strengths, potential shapes, and mass scales) are determined by the stabilization of fields at the GUT or Planck scale[Vilenkin1985].
Table 1 summarizes the hierarchy of these parameters. Below, we detail the specific high-energy mechanisms that constitute the coordinates of theory space and determine the specific constants we observe: Dimensionality, Fluxes, Moduli, and Symmetry Breaking.
2.1 Spacetime Structure and Algebra
We treat the Planck-scale coherent state as the initial boundary condition for the effective theory. We remain agnostic regarding the mechanism that populated this state–whether via tunneling from nothing (nucleation), a quantum bounce, or an eternal quantum static state. Our focus is exclusively on the subsequent dynamical selection of physical laws via decoherence as the system evolves away from this highly symmetric limit.
The most fundamental constants are those that define the algebraic structure of the quantum spacetime manifold itself. In our framework, these are treated as the rigid scaffolding of the theory space, often set to unity in natural units, but physically distinct in their dimensional roles.
  • Constants: Speed of light ($c$), Planck constant ($h$), Boltzmann constant ($k$), Stefan-Boltzmann constant ($\sigma$).
  • Mechanism: These are fixed by the selection of the fundamental operator algebra and the macroscopic dimensionality $D$ of the uncompactified spacetime.
  • Physical Role: The speed of light $c$ defines the causal structure (Lorentz invariance) of the manifold, while $h$ sets the scale of quantum commutation relations.
2.2 The First Transition: Dimensional Compactification
Before the universe can select specific interaction strengths or particle masses, it must establish the dimensionality of spacetime itself. In theories attempting to unify gravity with the standard model, such as string theory, the universe begins with a total dimensionality $D_{\text{total}}$ (e.g., 10 or 11). The first and most fundamental phase transition is Compactification, where the spatial dimensions split into two distinct sets:
$D_{\text{total}}\longrightarrow D_{\text{macro}}+d_{\text{micro}}$
Here, $D_{\text{macro}}$ represents the extended, observable spatial dimensions (typically 3), while $d_{\text{micro}}$ denotes the hidden, curled-up dimensions.
The size and shape of these extra dimensions are not fixed constants but are determined by the vacuum expectation value of a scalar field known as the radial modulus (or volume modulus). In the early high-energy “foam,” this modulus is a fluctuating quantum variable. If the modulus fluctuates, the fundamental couplings (such as the gravitational constant $G$ and the fine-structure constant $\alpha$) fluctuate with it. Consequently, the freezing of the radial modulus during the Compactification era is the prerequisite event that stabilizes the laws of physics. Only after the modulus settles into a minimum of its potential can the effective 4D field theory emerge.
2.3 Flux Integers
Flux integers are discrete quantum numbers ($N_{i}\in\mathbb{Z}$) arising from generalized magnetic fields threading the topological cycles (loops or holes) of the compactified extra dimensions.
  • Constants: Vacuum Energy (Cosmological Constant $\Lambda$).
  • Mechanism: Fluxes provide the “quantized steps” that generate the potential landscape. They do not set continuous couplings directly but determine which valley of the potential the universe can settle into.
  • Physical Role: The sum of flux energies and non-perturbative effects generates the vacuum energy density $\Lambda$. The discreteness of the flux is crucial for stabilizing the volume of the extra dimensions.
2.4 Moduli Fields: Couplings and Mass Blueprints
This is the critical section for the “fine-tuned” parameters. Moduli are continuous scalar fields ($\chi$) that parameterize the geometric deformations (radius, shape) of the extra dimensions. When these fields stabilize at the GUT scale (forming a condensate $\langle\chi\rangle$), they freeze the strength of all forces and the “blueprint” for future mass generation.
  • Gravity ($G$): The gravitational constant is determined by the Volume Modulus. In string theory, the strength of gravity in 4D is inversely proportional to the volume of the compactified dimensions ($G_{4D}\propto G_{fund}/V_{compact}$).
  • Electromagnetism ($\alpha,e,\epsilon_{0},\mu_{0}$): The fine-structure constant $\alpha=\frac{e^{2}}{4\pi\epsilon_{0}\hbar c}$ is set by the Dilaton VEV. Once $\alpha$ is fixed, the elementary charge $e$ and the vacuum response parameters $\epsilon_{0}$ and $\mu_{0}$ are determined.
  • Composite Constants ($\Phi_{0},G_{0},F$): Derived quantum constants such as the Magnetic Flux Quantum ($\Phi_{0}=h/2e$), Conductance Quantum ($G_{0}=2e^{2}/h$), and Faraday Constant ($F$) are locked in once the fundamental charge $e$ and Planck constant $h$ are set.
  • Nuclear Mass ($m_{p}$): The proton mass is determined by the QCD confinement scale $\Lambda_{QCD}$. This scale is set by the value of the strong coupling constant $\alpha_{s}$ at the GUT scale running down via the Renormalization Group. Thus, $m_{p}$ is effectively fixed at the GUT scale by the modulus determining $\alpha_{s}$.
  • Elementary Mass ($m_{e},R_{\infty}$): The electron mass is the product of the Higgs VEV ($v$) and the electron Yukawa coupling ($y_{e}$). Both $v$ (determined by parameters in the Higgs potential) and $y_{e}$ (determined by intersection geometries in string theory) are functions of the moduli. Therefore, the instruction set for the electron mass and the Rydberg constant ($R_{\infty}$) is frozen pre-inflation, ensuring no domain walls form during the later Electroweak transition.
2.5 Symmetry Breaking and Vacua
While the previous mechanisms operate at the Planck or String scale, the final form of the laws of physics involves the breaking of high-energy symmetries (e.g., $SU(5)$ or $SO(10)$) down to the Standard Model group.
  • Constants: Particle generations, Mixing angles (CKM/PMNS matrices).
  • Mechanism: Discrete choices, such as the presence of Wilson lines or discrete symmetries (like Parity), differentiate between vacua that share the same continuous moduli values. These choices determine the “texture” of physics, such as the number of particle families.
Setting Era Mechanism Constants Determineda Key Reference
String Scale
Stabilization of Radial Moduli;
Topology Change.
$G$ (Gravitational Const.)
$D$ (Dimensions)
Candelas et al. (1985) [Candelas1985]
Planck Era
Euclidean Path Integral;
Geometric Flux Cancellation.
$\Lambda$ (Vacuum Energy) Baum (1983) [Baum1983];
Hawking (1984) [Hawking1984]
GUT Scale
(Symmetry Breaking)
Flux Quantization;
Moduli VEVs (Dilaton);
Group Theory Factors.
$\alpha,\alpha_{s},e,\mu_{0},\epsilon_{0}$
$m_{e},m_{p},m_{n}$
(via Yukawas & $\Lambda_{\text{QCD}}$ seed)
Kaplunovsky (1988) [Kaplunovsky1988];
Susskind (2003) [Susskind2003]
  • a
    While masses ($m_{e},m_{p}$) manifest at low energy, their values are strictly determined by the dimensionless couplings (Yukawas and $\alpha_{s}$) fixed at the pre-inflationary GUT scale.
Table 1: The hierarchy of constants fixed in the pre-inflationary universe. All fundamental parameters are determined by geometric and topological transitions occurring before the onset of cosmic inflation, ensuring their universality across the observable cosmos.
3 Enlarged Configuration Space
3.1 Superspace and Theory Space
Canonical quantum gravity is traditionally defined over superspace ($\mathcal{S}$), the space of all possible 3-geometries $h_{ij}]$ and matter fields $\phi$ on a spatial slice $\Sigma$. The Wheeler–DeWitt equation, $\hat{\mathcal{H}}\Psi=0$, describes the static quantum state of the universe over this space. However, this equation requires the constants of nature to be inserted as fixed numbers.
To describe a quantum origin for the laws of physics, we must treat these constants as coordinates. We introduce theory space ($\mathcal{T}$), a manifold where each point $\mathcal{T}=(D,X)$ represents a complete, self-consistent set of physical laws derived from a fundamental theory with total dimensionality $D_{\text{fund}}$.
The coordinates are defined as follows:
  • Macroscopic Dimension ($D$): The integer number of large, expanding spatial dimensions. This acts as a discrete sector label.
  • Internal Parameters ($X$): A vector containing all continuous and discrete variables required to define the full $D$-dimensional effective Lagrangian and vacuum structure of the resulting U-sector. This includes the moduli and flux parameters associated with the stabilization of the $d=D_{\text{fund}}-D$ compactified dimensions.
The full configuration space is the union of superspace and theory space:
$\widetilde{\mathcal{Q}}=\mathcal{S}\times\mathcal{T}=\{[h_{ij}],\phi,D,X\}.$ (1)
In this notation, $D$ selects the specific effective field theory (the dimensionality of the superspace fiber), while $X$ determines the parameters (masses and couplings) within that theory.
4 Quantum Structure of Constants
4.1 U-Sector Hilbert Spaces and the Direct-Sum Structure
Definition.
We define a U-sector not merely as a distinct quantum state, but as a distinct Hilbert space. It is insufficient to treat universes with different physical constants simply as orthogonal vectors within a single, shared Hilbert space. Because the fundamental constants (such as the electron mass or fine-structure constant) appear in the definition of the canonical commutation relations and the Hamiltonian operator itself, states associated with different constants belong to inequivalent representations of the field algebra.
Therefore, the total configuration space cannot be modeled as a tensor product of a geometry space and a parameter space. Instead, the appropriate mathematical structure is a direct sum of the Hilbert spaces associated with every possible choice of physical laws[Roberts1990].
Let $\mathcal{H}^{(D,X)}$ denote the standard Hilbert space of canonical quantum gravity for a fixed background theory $\mathcal{T}=(D,X)$ (i.e., the space of wavefunctions $\Psi[h_{ij},\phi]$ governed by a specific Hamiltonian constraint). The total, or “Grand,” Hilbert space is:
$\mathcal{H}_{\rm grand}=\bigoplus_{(D,X)\in\mathcal{T}}\mathcal{H}^{(D,X)}.$ (2)
In this framework, the meta-wavefunctional of the universe is a vector with components in each sector:
$\Psi=\bigl(\Psi_{D_{1},X_{1}},\,\Psi_{D_{2},X_{2}},\,\ldots\bigr),\qquad\Psi_{D_{i},X_{i}}\in\mathcal{H}^{(D_{i},X_{i})}.$ (3)
with the inner product
$\langle\Psi,\Psi\rangle=\sum_{(D,X)}\langle\Psi_{D,X},\,\Psi_{D,X}\rangle=1.$ (4)
Unlike a standard superposition, the components $\Psi_{D_{i},X_{i}}$ evolve under different Hamiltonians $\hat{\mathcal{H}}_{D_{i},X_{i}}$. Consequently, there are no local physical operators that can map states from one sector to another, rendering them physically disjoint (superselected).
Ontological Note.
We acknowledge that we have no way of knowing how nature physically “stores” or operates on the wavefunction of the universe. The direct-sum formalism presented here is a mathematical suggestion–a rigorous way to describe a state space where the laws of physics themselves are quantum variables, while respecting the fact that observers within the system perceive those laws as fixed classical backgrounds.
4.2 The Extended Path Integral
It is plausible that Nature admits a single underlying, fully self-consistent microscopic theory. Even a unique fundamental theory, however, may possess a vast set of metastable vacua, compactification branches, and flux choices. In this picture, the multiplicity of possible constants does not reflect a multiplicity of unrelated “theories” but rather the many high-energy realizations of one underlying theory.
To formalize this, we extend the gravitational path integral to include the theory space variables $\mathcal{T}\equiv(D,X)$ on the same footing as the geometric and matter degrees of freedom. Wheeler’s original spacetime foam concerned fluctuations in geometry at every point in space. We extend this to a global fundamental foam, where the constants themselves fluctuate. The extended functional integral becomes:
$\mathcal{Z}_{\rm ext}\;=\;\int\mathcal{D}_{\mathcal{T}}\mathcal{D}_{g}\mathcal{D}_{\phi}\;e^{\left\{i\,S[g,\phi,\mathcal{T}]\right\}}.$ (5)
The action $S[g,\phi,T]$ governing a specific sector $T=(D,X)$ is given by the integral over the $D$-dimensional spacetime manifold $\mathcal{M}_{D}$:
$S[g,\phi,T]\;=\;\int_{\mathcal{M}_{D}}d^{D}x\sqrt{-g}\left[\frac{R-2\Lambda(X)}{16\pi G(X)}+\mathcal{L}_{\rm matter}\big(\phi,\nabla\phi;g_{i}(X)\big)\right].$ (6)
Here, the fundamental parameters—such as the effective gravitational constant $G(X)$, the cosmological constant $\Lambda(X)$, and the gauge couplings $g_{i}(X)$ within the matter Lagrangian—are determined by the moduli $X$.
Here, the fundamental parameters–such as the effective gravitational constant $G(X)$, the cosmological constant $\Lambda(X)$, and the gauge couplings $g_{i}(X)$ within the matter Lagrangian–are not constants but functions of the theory-space coordinates $X$. Here, the integration measure is nested: the global integration $\mathcal{D}_{\mathcal{T}}$ sums over all possible U-sectors (the choice of laws), while the inner functional integrations $\mathcal{D}_{g}\mathcal{D}_{\phi}$ sum over all local histories within that specific sector. $\mathcal{Z}_{\rm ext}$ thus represents the total amplitude of the cosmic landscape.
A crucial advantage of the path integral formulation is that it provides a natural definition for the probability measure on theory space, distinct from the Anthropic Principle.
$P(\mathcal{T}_{i})\;=\;\frac{\left|Z(\mathcal{T}_{i})\right|^{2}}{\int_{\mathcal{T}}\mathcal{D}_{\mathcal{T}}\left|Z(\mathcal{T})\right|^{2}},$ (7)
where $Z(T_{i})$ is the path integral over geometries and matter fields for the specific fixed laws $T_{i}$:
$Z(\mathcal{T}_{i})=\int\mathcal{D}_{g}\mathcal{D}_{\phi}\;e^{\,i\,S[g,\phi,\mathcal{T}_{i}]}.$ (8)
In this expression, the denominator serves as the normalization factor, summing the quantum weights of all possible U-sectors in the landscape. This form of the action, where couplings are promoted to functions of scalar moduli, is the standard low-energy effective action found in string theory [Polchinski1998] and scalar-tensor theories of gravity [Uzan2011]. $\mathcal{Z}_{\rm ext}$ implies that the relative probability of a specific U-sector $\mathcal{T}_{i}$ is weighted not just by the action $S$, but by the volume of the configuration space $\mathcal{D}_{T}$ surrounding it. This suggests an entropic selection principle: U-sectors that occupy larger volumes in the landscape of moduli and fluxes are quantum-mechanically favored. Thus, the observed constants are likely not just anthropically valid, but also correspond to “broad” minima in the theory potential $V(\chi)$, rather than narrow, fine-tuned clefts.
5 The Meta-Wheeler–DeWitt Equation
We now translate the extended path integral into the canonical formalism. We begin by reviewing the standard constraint equation to highlight where the constants enter as fixed constraints, before generalizing them to dynamical operators.
5.1 The Standard Wheeler–DeWitt Constraint
In standard canonical quantum gravity, the dynamics of the universe are governed by the Wheeler–DeWitt equation [DeWitt1967]. This equation arises from the quantization of the Hamiltonian constraint of General Relativity. For a wavefunction $\Psi[h_{ij},\phi]$ defined on the superspace of 3-geometries $h_{ij}$ and matter fields $\phi$, the constraint is:
$\hat{\mathcal{H}}\Psi=\left(\hat{\mathcal{H}}_{\rm grav}+\hat{\mathcal{H}}_{\rm matter}\right)\Psi=0.$ (9)
Explicitly, the gravitational part (in the position representation) typically takes the form:
$\hat{\mathcal{H}}_{\rm grav}=-16\pi G\cdot G_{ijkl}\frac{\delta^{2}}{\delta h_{ij}\delta h_{kl}}-\frac{\sqrt{h}}{16\pi G}\left({}^{(3)}R-2\Lambda\right).$ (10)
Crucially, in this standard formulation, the parameters $G$, $\Lambda$, and the coupling constants inside $\hat{\mathcal{H}}_{\rm matter}$ appear as fixed c-numbers. They are external inputs to the theory.
5.2 Generalization to Theory Space
We propose that the fundamental constants are not fixed parameters but eigenvalues of field operators acting on the theory-space coordinates $(D,X)$. We promote the Hamiltonian constraint to a “Meta-Wheeler–DeWitt” operator:
[…]
Here, $\hat{\mathcal{H}}_{\rm grav}$ and $\hat{\mathcal{H}}_{\rm matter}$ retain their standard forms, but their coefficients (the constants) are now operators dependent on the theory-space variables.
The third term, $\hat{\mathcal{H}}_{\rm theory}$, is new. It governs the dynamics of the constants themselves—the transitions between different values of $D$ and $X$. It typically takes the form:
$\hat{\mathcal{H}}_{\rm theory}=\hat{T}_{\rm smooth}+\hat{T}_{\rm discrete}+\hat{U}_{\rm theory}(D,X).$ (12)
$\hat{T}_{\rm smooth}$ describes the evolution of continuous moduli, while $\hat{T}_{\rm discrete}$ allows for tunneling between discrete sectors (fluxes and dimensions).
6 Sector Separation
We argue that the orthogonality of U-sectors is dynamically established rather than imposed as a fundamental law. The claims regarding non-interference are emergent properties of the low-energy physics.
6.1 Emergent Superselection
The tension between the pre-inflationary fully coupled Hamiltonian and the isolated nature of physical laws is resolved by cosmic expansion. The total Hamiltonian transitions from a coupled state to a block-diagonal state due to the dynamic suppression of the tunneling operator:
  • Primordial Coherence (Coupled): During the Planck era, the discrete kinetic term $\hat{T}_{\rm discrete}$ inside $\hat{\mathcal{H}}_{\rm theory}$ is non-vanishing (tunneling amplitudes $\kappa\neq 0$). This allows the wavefunction to spread across different sectors.
  • Dynamic Superselection (Separated): As the universe expands, the tunneling amplitudes $\kappa$ become exponentially suppressed ($\Gamma\sim e^{-\frac{S_{instanton}}{\hbar}}\to 0$). The off-diagonal terms $\hat{T}_{ij}$ vanish. The Hamiltonian constraint $\hat{\mathcal{H}}_{\rm meta}$ transitions to a stable block-diagonal form:
    $\hat{\mathcal{H}}_{\rm meta}\xrightarrow{t\to\infty}\bigoplus_{(D,X)}\left(\hat{\mathcal{H}}_{\rm grav}^{(D,X)}+\hat{\mathcal{H}}_{\rm matter}^{(D,X)}\right).$ (13)
Thus, for all observers in the post-Planck classical regime, the sectors are rendered dynamically and effectively superselected. No local operator can bridge the gap between the now-orthogonal Hilbert spaces.
6.2 Localization and the Parent Observable
In our localized U-sector, the expectation value of $\chi$ ($\chi_{*}$) fixes the value of parameters like the Higgs VEV ($v$) and the strong coupling constant ($\alpha_{s}$). These are all directly observable quantities in particle physics experiments:
$\chi_{*}\longrightarrow\text{Higgs VEV }(v)\longrightarrow\text{Electron Mass }(m_{e}).$
$\chi$ acts as the “Parent Observable”: if one accepts that $m_{e}$ is an observable, and $m_{e}$ is fixed by the classical value of $\chi$, then $\chi$ must be the quantum source of that observable.
The absence of interference is a structural feature of the Hilbert space that emerges after the Planck era, not a result of environmental scattering. The only residual role of the initial superposition is to fix the relative weights of the decohered U-sectors.
7 Cosmological Evolution of the Meta-Wavefunction
7.1 The Emergence of the Standard Model Hamiltonian
We can derive the effective physics observed today by expanding the fundamental “Eternal” Hamiltonian around the stabilized vacuum $\chi_{*}$. We decompose the total Hamiltonian into the dynamics of the constants ($\chi$), the matter fields ($\phi$), and gravity ($g$):
$\hat{\mathcal{H}}_{\rm eternal}=\underbrace{\hat{K}_{\chi}+V(\chi)}_{\text{Moduli Sector}}+\underbrace{\hat{K}_{\rm grav}+\hat{V}_{\rm grav}(\chi)}_{\text{Gravity Sector}}+\underbrace{\hat{K}_{\rm matter}+\hat{V}_{\rm matter}(\phi,\chi)}_{\text{Matter Sector}}.$ (14)
Here, the dependencies on $\chi$ are explicit:
  • $\hat{V}_{\rm grav}$ depends on $\chi$ through the gravitational constant $G(\chi)$.
  • $\hat{V}_{\rm matter}$ depends on $\chi$ through the mass terms $m(\chi)$ and couplings $\lambda(\chi)$.
  • $\hat{K}_{\chi}$ represents the “velocity” of the constants, suppressed by the effective mass $M_{\rm eff}\propto a^{3}$.
Expanding the field $\chi=\chi_{*}+\delta\chi$, we recover the Standard Model plus suppressed interaction terms:
$\hat{\mathcal{H}}_{\rm eternal}\approx\underbrace{\left[\hat{K}_{\rm grav}+\hat{V}_{\rm grav}(\chi_{*})+\hat{K}_{\rm matter}+\hat{V}_{\rm matter}(\phi,\chi_{*})\right]}_{\textbf{Standard Model (Fixed Laws)}}+\underbrace{\left[\frac{\hat{P}_{\delta\chi}^{2}}{2M_{\rm eff}}+\frac{1}{2}V^{\prime\prime}(\chi_{*})(\delta\chi)^{2}\right]}_{\textbf{Frozen Moduli}}+\hat{\mathcal{H}}_{\rm int}.$ (15)
The term $\hat{\mathcal{H}}_{\rm int}$ represents the coupling between the fluctuations of constants and the observable physics:
$\hat{\mathcal{H}}_{\rm int}=\delta\chi\left(\frac{\partial\hat{V}_{\rm grav}}{\partial\chi}\Bigg|_{\chi_{*}}+\frac{\partial\hat{V}_{\rm matter}}{\partial\chi}\Bigg|_{\chi_{*}}\right).$ (16)
The first term in parentheses represents a time-varying gravitational constant ($\dot{G}$), and the second represents time-varying particle masses ($\dot{m}$). Today, we ignore $\hat{\mathcal{H}}_{\rm int}$ because the massive potential $V^{\prime\prime}(\chi_{*})$ locks $\delta\chi\to 0$, rendering these variations experimentally undetectable.
7.2 Backward Evolution: The Present as Boundary Condition
To determine the initial state of the universe, we do not rely on speculative high-energy postulates. Instead, we adopt a rigorous empirical approach: we treat the present observed state as the fixed boundary condition and integrate the quantum dynamics backward to the Planck era.
Let the current state of the universe at scale factor $a_{\rm now}$ be defined by the observed values of the constants $X_{\rm obs}$. In the theory space representation, the wavefunction today is effectively a delta function (or a very sharp Gaussian) peaked at the observed modulus value $\chi_{*}$:
$\Psi(a_{\rm now},\chi)\;\approx\;\delta(\chi-\chi_{*})\otimes\Psi_{\rm geom}(a_{\rm now}).$ (17)
This represents a universe with classically fixed laws of physics.
We then solve the Wheeler–DeWitt equation in the direction of decreasing $a$. The evolution is governed by the $a$-dependent effective mass of the theory-space variables, which scales as $M_{\rm eff}\propto a^{3}V^{\prime\prime}(\chi)$.
  • The Classical Era ($a\gg l_{\rm Pl}$): For macroscopic $a$, $M_{\rm eff}$ is large. The wavefunction remains tightly localized around $\chi_{*}$. The constants appear immutable.
  • The Quantum Era ($a\to 0$): As we approach the Planck scale, the effective mass $M_{\rm eff}$ vanishes. The restoring force that pins $\chi$ to $\chi_{*}$ disappears. Due to the uncertainty principle, a state that is sharply localized at late times must disperse as the system evolves backward into the low-mass regime.
Consequently, the backward integration demonstrates that the unique, sharp state we observe today originates from a coherent superposition over the landscape potential at the Planck scale. The asymptotic flatness of the initial state is not an assumption; it is a mathematical necessity required to produce the sharp classical state we observe today.
Mathematically, this flatness arises directly from the structure of the Hamiltonian constraint in the limit $a\to 0$. As the scale factor vanishes, the term governing the curvature of the wavefunction with respect to $\chi$ dominates. The constraint equation for the theory-space component $\xi(\chi)$ effectively becomes:
$\frac{1}{2M_{\rm eff}(a)}\frac{\partial^{2}\xi}{\partial\chi^{2}}\approx 0\quad\implies\quad\frac{\partial^{2}\xi}{\partial\chi^{2}}\approx 0.$ (18)
Since the effective mass $M_{\rm eff}$ vanishes, the kinetic cost of spatial gradients in theory space disappears. The solution to $\partial_{\chi}^{2}\xi=0$ that satisfies the no-boundary proposal (regularity) is a constant, $\xi(\chi)\sim\text{const}$. Thus, the derivative vanishes, $\frac{\partial\Psi}{\partial\chi}\to 0$, indicating a state of maximum symmetry where no value of the parameter $\chi$ is favored over another.
7.3 Forward Evolution: Entanglement and Branching
We now consider the forward evolution from the initial state. We assume the universe begins in a state of Primordial Coherence, where $\Psi(a\approx 0)$ is a broad superposition over the theory space $\mathcal{T}$.
$\Psi(a\approx 0)\sim\int_{\mathcal{T}}d\chi\,c(\chi)\,|a=0,\chi\rangle\otimes|\Omega_{\phi}\rangle$ (19)
Here, $\chi$ represents the continuous coordinates of theory space (moduli), and $|\Omega_{\phi}\rangle$ is the vacuum state of the matter fields.
As the universe expands ($a$ increases), the interaction term $V(\phi,\chi)$ in the Hamiltonian creates entanglement between the theory parameters and the matter dynamics.
$\Psi(a)\longrightarrow\int_{\mathcal{T}}d\chi\,c(\chi)\,\xi(\chi;a)\,\psi_{\chi}(a,\phi)$ (20)
The single wavefunction splits into an infinite superposition of branches. In each branch $\chi$, the matter field $\phi$ evolves according to a unique effective Hamiltonian $\hat{H}_{\chi}$. For example, if $\phi$ is the inflaton, different $\chi$ branches corresponds to universes with different inflationary potentials, leading to different durations of inflation and different primordial power spectra.
7.4 Sector Selection via Freezing
The critical transition occurs when the scale factor $a$ crosses the threshold for moduli stabilization. The “force” of the potential $V(\chi)$ grows, and the broad distribution $\xi(\chi;a)$ fragments into localized wavepackets centered on the local minima $\chi_{i}$ of the landscape:
$\Psi(a\to\infty)\approx\sum_{i}C_{i}\,\Psi_{\text{U-sector }i}(a,\phi).$ (21)
This process is the quantum origin of the U-sectors. The continuous initial state is projected onto the discrete basis of stable vacua. The relative probability of our universe is determined by the weight $|C_{*}|^{2}$ of the wavepacket that settles into our specific minimum $\chi_{*}$.
7.5 Comparison with Standard Boundary Conditions
The boundary condition proposed here, a delocalized “flat” state in theory space, differs fundamentally from standard approaches to the initial state of the universe. In traditional quantum cosmology, boundary proposals such as the Hartle-Hawking (No-Boundary) state or the Vilenkin (Tunneling) state are primarily concerned with the geometric degree of freedom $a$.
  • Standard Approach: In models with a scalar field $\phi$, the wavefunction is typically assumed to be the ground state of the matter Hamiltonian, $\Psi\sim e^{-\phi^{2}}$. This assumes the effective potential $V(\phi)$ is already fixed and steep enough to localize the field near the origin. It implicitly assumes the laws of physics (the parameters of $V$) are pre-determined.
  • Theory-Space Approach: In our framework, the potential $V(\chi)$ governing the constants vanishes or becomes negligible at the Planck scale due to the vanishing effective mass. Consequently, the correct boundary condition in the $\chi$-direction is not a localized Gaussian, but a state of maximum entropy:
    $\lim_{a\to 0}\Psi(\chi)\sim\text{const}.$ (22)
    This flatness implies that no specific U-sector is preferred ab initio. The selection of physical laws is not an initial condition imposed at the Big Bang, but a dynamical outcome of cosmic history. The universe begins in a superposition of all possible theories, and only the expansion of spacetime breaks this symmetry, collapsing the flat distribution into the sharp peaks we observe today.
8 Discussion
8.1 The Peak Prediction: A Testable Constraint
Although the configuration space that allows complexity ($\Omega_{\rm obs}\subset\mathcal{T}$) is small relative to the huge space of possibilities, it occupies a significant volume. The quantum weighting framework predicts that the observed values of physical constants is most likely to sit near the maximum of the probability distribution, $P(\mathcal{T})$, within this habitable zone. This is the $\max P(\mathcal{T})$ prediction. If the observed universe were found in a low-probability tail region of $\Omega_{\rm obs}$, it would suggest the existence of a secondary, unknown filter beyond the initial quantum weighting. We are currently limited from testing this prediction because the volume and structure of the complexity-supporting region $\Omega_{\rm obs}$ have not been fully mapped; however, future computational work may render this constraint testable.
8.2 A Criterion for Fundamental Theories
Beyond explaining fine-tuning, the extended path integral formulation offers a quantitative tool for comparing candidate microscopic theories. In Bayesian inference, the relative plausibility of two models, $\mathcal{M}_{A}$ and $\mathcal{M}_{B}$, given observational data $D_{\rm obs}$ (i.e., the measured constants), is determined by the ratio of their evidences (the Bayes Factor):
$\frac{P(\mathcal{M}_{A}|D_{\rm obs})}{P(\mathcal{M}_{B}|D_{\rm obs})}=\frac{P(D_{\rm obs}|\mathcal{M}_{A})}{P(D_{\rm obs}|\mathcal{M}_{B})}\cdot\frac{P(\mathcal{M}_{A})}{P(\mathcal{M}_{B})}.$ (23)
In our framework, the “evidence” $P(D_{\rm obs}|\mathcal{M})$ corresponds to the total quantum measure of the U-sectors that are compatible with observation. If $\Omega_{\rm obs}\subset\mathcal{T}$ is the region of theory space containing constants compatible with our existence, then the evidence provided by a specific fundamental theory is:
$E(\mathcal{M})\;=\;\int_{\Omega_{\rm obs}}\mathcal{D}_{T}^{\mathcal{M}}\left|Z(T)\right|^{2}.$ (24)
This functional allows us to rate competing hypotheses about high-energy physics (e.g., different compactification manifolds or flux stabilization mechanisms). A theory that naturally generates a “fertile” landscape–one where $\Omega_{\rm obs}$ occupies a large volume of the configuration space–will yield a larger $E(\mathcal{M})$ and is statistically favored over a theory that requires extreme fine-tuning to produce a single viable U-sector.
Thus, $\mathcal{Z}_{\rm ext}$ transforms the measure problem from an interpretational difficulty into a rigorous selection criterion for ultraviolet completions of the Standard Model.
8.3 Causality and the Horizon Problem
One might object, following Bousso and Susskind [BoussoSusskind2012], that light-speed causality forbids a “global” decoherence event across a spatially infinite universe, as no causal signal can synchronize the symmetry breaking. However, our framework operates within the minisuperspace approximation, which describes the evolution of a homogeneous patch. Physically, this corresponds to the pre-inflationary era where the causal horizon was microscopic. The selection of a specific U-sector (a value for $\chi_{*}$) occurs within this initial causal patch. The subsequent period of cosmic inflation expands this single, homogeneous domain to macroscopic scales, pushing any domain walls far beyond the current observable horizon. Thus, for all observational purposes, the constants appear global and fixed.
8.4 A Prediction of a Multiverse Framework
A central challenge for any multiverse framework is the apparent lack of causal contact between sectors. Since U-sectors are separated by superselection rules, we have no known local operator that can interact with or retrieve information from a parallel theory branch.
However, this framework makes a specific, falsifiable prediction regarding the future of theoretical physics: there will never be a successful derivation of the Standard Model parameters from first principles.
If the constants of nature are quantum variables subject to early-universe branching, then their specific values are environmental accidents frozen by decoherence, not mathematical inevitabilities. Consequently, any search for a “Theory of Everything” that attempts to uniquely constrain the values of the fine-structure constant, electron mass, or cosmological constant from pure geometry or algebra is destined to fail. The prediction of this framework is that the fundamental theory will yield a probability distribution over these parameters, not a single point.
8.5 Relation to Previous Work
The conceptual unification of the cosmological multiverse with the Many-Worlds Interpretation has been proposed previously, notably by Nomura [Nomura2011] and Susskind/Bousso [BoussoSusskind2012]. However, these approaches typically rely on the semiclassical background of eternal inflation and bubble dynamics.
This work distinguishes itself by providing a fully canonical realization of this unification. Instead of relying on bubble nucleation, we derive the splitting of physical laws directly from the decoherence of the universal wavefunction in the extended configuration space. By utilizing the Wheeler-DeWitt formalism, we show how “theory selection” emerges as a stationary feature of the quantum state $\Psi$, independent of the specific temporal slicing issues associated with the measure problem in eternal inflation.
9 Conclusion
We have developed an extension of canonical quantum cosmology in which the constants of nature are promoted to quantum variables. By enlarging the Wheeler–DeWitt configuration space to include the theory-space coordinates $(D,X)$, the universal wavefunction gains support not only over classical geometries but also over possible low-energy laws themselves. Each assignment $(D,X)$ defines a U-sector with its own Hamiltonian constraint, and the grand quantum state is a superposition over these sectors, thereby extending the Everett Many Worlds interpretation to include many worlds of different physical laws.
Specifically, we have accomplished the following:
  1. 1.
    We defined the Grand Hilbert Space as the direct sum of U-sectors, each representing a distinct set of physical laws (Hamiltonian constraint).
  2. 2.
    We derived the Meta-Wheeler–DeWitt equation, which governs the dynamics of the theory parameters. We showed that in the Planck era, the kinetic terms for these parameters are active, creating a state of primordial coherence where the laws of physics effectively fluctuate.
  3. 3.
    We identified the mechanism of Sector Separation. As the universe expands, the effective mass of the theory moduli grows, freezing them into fixed values. The source of decoherence in the early universe is the interaction between these global parameters and the inhomogeneous degrees of freedom (matter and metric fluctuations), which renders the branches of different U-sectors physically disjoint.
In this framework, the fine-tuning problem is reinterpreted as a quantum weighting problem: the observed constants correspond to the sector(s) for which the meta-wavefunction has non-negligible amplitude. While the vast majority of the cosmic wavefunction may describe sterile vacua, we necessarily inhabit one of the rare branches compatible with complexity.
Beyond its interpretational value, this formalism offers a quantitative tool for high-energy physics. We proposed that the total integrated amplitude over theory space, $\mathcal{Z}_{\rm ext}$, serves as a Bayesian evidence metric, allowing for the statistical comparison of rival microscopic theories (e.g., different string compactifications) based on their fertility in generating habitable sectors.
Finally, this framework makes a specific, falsifiable prediction if it is the correct one: there will never be a successful, purely mathematical derivation of the Standard Model parameters.** The ultimate theory will yield a probability distribution, not a point solution. This suggests that the laws of physics we observe are not fixed external constraints, but frozen quantum accidents—local fossils of the early universe’s decoherence history.